# Simultaneous FUV and NUV observations of T Tauri stars with UVIT/AstroSat: probing accretion process in young stars

### v1

@@ 2.3 Spectral energy distribution
Out of 11 UVIT-detected TTS candidates, only seven are detected in both FUV and NUV bands, while the others are detected only in the NUV bands. Therefore, we performed two different approaches to fit the observed SEDs to estimate the stellar parameters. We made a two-component fit to the UV and optical part of the SEDs of the seven TTSs detected in both FUV and NUV bands; we used a stellar photospheric model and a blackbody (which represents the excess over the photosphere emission due to accretion shocks) to fit the observed SEDs to determine the excess UV luminosity, its equivalent temperature, and stellar parameters of the stars. We did not include IR regions for the SED fit as the excess emission in the IR region comes from the disk. For the remaining four TTSs, we only fit the photospheric models of dwarf stars to the optical part of the observed SEDs to estimate stellar parameters.
To construct observed SEDs, we cross-matched our sample with Gaia-DR3 (Gaia Collaboration et al., 2021; Babusiaux et al., 2022), APASS DR9 (Henden et al., 2015), and PanSTARRS (Magnier et al., 2020) for the optical; and 2MASS (Skrutskie et al., 2006), Spitzer and WISE (Wright et al., 2010) for IR region. We also included GALEX GR6/7 data for this analysis. We have used the virtual observatory (VO) tool, VOSA (VO SED Analyzer; Bayo et al., 2008) to generate SEDs by converting magnitudes into flux values for corresponding filters and then to fit the observed flux distribution with theoretical model spectra. We have obtained the distances from the Gaia-DR3 (Bailer-Jones et al., 2021). VOSA performs multiple iterations to get best-fitted spectra to observed flux distribution by varying T, log(g), metallicity, extinction, and scaling factor () values and gives the best-fitted parameters after performing a minimization. As these TTSs are in the solar neighbourhood, we consider that they all have similar metallicity as the Sun. We kept extinction as a free parameter to estimate its value from the best-fitted spectra. The scaling factor () is used to scale the model flux to match the observed flux and is defined as where is the radius of the star and D is the distance to the star. For the stellar photosphere, we have used the BT-Settl-CIFIST model (Allard et al., 2011). We have used log(g) values between 4 to 5 and the full range for T from 1200 to 7000 K. In the case of the black body, we used a range of temperatures from 5000 to 20000 K.
@@
where N is the number of photometric data points, n is the number of input-free parameters, is the observed flux, and is the model flux.
@@
In Figure 2, we have shown SEDs of the seven TTSs (detected in both FUV and NUV) fitted with model spectra. The observed fluxes are shown as cyan and blue points, where cyan points fall in the UV and optical regions of the energy distribution which is included in the dual component fitting of model spectra. The blue points fall in the IR region and are not included in the fit. The black and grey lines represent the best-fitted black-body and BT-Settl-CIFIST spectra, respectively, to the cyan points. The combined flux of these two model spectra is shown as the red dashed line. The red points indicate the combined flux in different bands. The overlap between red and cyan points indicates the goodness of the two-component fitting. We notice that in the case of HD283782, combined flux matches very well in the IR region (not included in the fit) as well. Therefore, there is no IR excess present in HD283782, which indicates that the TTS does not have hot/warm dust around it. The values of reduced , temperatures, logg, and extinction (A) corresponding to the best-fitted spectra are noted in the legends. The names of the sources are mentioned at the top of each plot. The figure suggests that FN Tau, V836 Tau, and HD283782 have high values, which could be due to variability.
@@
In Figure 3, we have shown SEDs of TTSs with no FUV detections. Due to the absence of FUV points, only the optical region of the observed SED is fitted with synthetic dwarf spectra (BT-Settl-CIFIST) to get the stellar parameters of these TTSs. We have used T, , and A as free parameters, and metallicity to be fixed at Solar value. Blue and cyan points in Figure 3 are observed SEDs, where only cyan points are included in the fitting process. The gray spectrum indicates the best-fitted model spectrum to the optical region of the observed SED. The red point indicated the expected flux due to the best-fitted model spectrum. The best-fitted parameters are mentioned in the legend. Anon 1 shows almost no excess in the IR regions, indicating a lack of hot/warm dust around it.
@@ 3 Results and Discussion
The SED analysis provides us with fundamental parameters of the TTSs, i.e., T, logg, radius, mass, and bolometric luminosity (L) of the central stars. As mentioned on the VOSA website, the fitting process and the predicted flux are relatively less sensitive to . This poses a challenge to put direct constraints on the masses of the components since a slight change in can create a large difference in the estimated mass. Therefore, we do not include the estimations of mass and in our analysis. The values of other parameters are listed in Table 3. The luminosity of the black-body provides a direct measure of the accretion luminosity/chromospheric luminosity. In Table 3, we listed the value of bolometric luminosity enclosed within the black-body spectrum (L) and black-body temperature (T). We have also estimated the spectral type of each TTS using T-spectral type relation (Pecaut & Mamajek, 2013) and listed in column 7 of Table 3. We find that TTS have a T range from 3000 K to 4600 K and spectral types of K and M. The T ranges from 6800 K to 10300 K.
3.1 UV excess as a indicator of accretion
@@
We notice in Figure 2 that V773 Tau and HD283782 emit negligible amounts of excess emission in NUV bands above the photospheric emission compared to that found in the FUV bands. Whereas, in the other five sources we observe a large excess emission in both NUV and FUV bands. In Figure 3, we further notice that Anon1 has less excess NUV bands compared to other sources. The presence/absence of NUV excess appear to divide sources into two different categories.
@@
However, we need to quantitatively estimate the amount of excess emission in all the sources and compare them before giving any conclusion. To quantify the excess flux observed in the FUV and NUV regions, we define excess emission in a particular band as . To demonstrate UV excess, we used the N245M filter in the NUV and the F148W band in the FUV regions, as these are common bands for all the sources. The difference in photospheric temperature or spectral types of the stars could also lead to variation in the excess emission. For a hotter star, the excess emission may appear low, which can also affect the blackbody temperature. Therefore, we have plotted NUV and FUV excess as a function of effective temperature (T) in the upper panel of Figure 4 and as a function of blackbody temperature (T) in the lower panel. T and T are estimated from SED fitting. The sources detected in both FUV and NUV are marked as blue points, while the sources with only NUV detection are marked in orange. We can see that there are two groups of population, independent of their spectral types and blackbody temperature, separated by a black dashed line. The presence of high UV excess (10 in NUV and 10 in FUV) suggests that these eight sources (FM Tau, CW Tau, FN Tau, BS Tau, V836 Tau, CIDA1, FO Tau and J0414) could possibly be CTTS, while V773 Tau, HD283782 and Anon1 are probable WTTS candidates with comparatively little excess in UV. Comparatively low UV excess with almost no IR excess in HD283782 indicates that it is an example of disk-less WTTS. A more careful inspection suggests that stars with higher UV excess (or the probable CTTS) also tend to have hotter T compared to the stars with lesser UV excess, with FN tau as an exception. However, we need more sources with simultaneous UV observations to confirm this aspect. We classify the stars emitting 10 (10) excess emission in NUV (FUV) as CTTS, while the other group of stars are classified as WTTS. Our classification based on the UV excess is listed in the Table 1. We, however, require further evidence from spectroscopic observations before confirming their classification as CTTS or WTTS, as CTTSs are expected to show strong H emission compared to WTTS.
| TTS | T | T | Radius | L | Spectral Type | L | ||
| Name | (K) | (K) | (R) | (L) | (L) | (10 Myr) | ||
@@ 3.2 H line emission as an accretion indicator
We analyzed spectra obtained from both the HFOSC/HCT and LAMOST for our sources. We estimated the equivalent width of H line emission (EW(H)) after subtracting continuum flux from the spectra. In Table 2, we have listed the EW(H) for each source obtained from the multi-epoch observations with different telescopes (column 3). Names of the telescopes and the dates of observations are also mentioned in columns 2 and 3, respectively. We notice that EW(H) of FM Tau and V773 Tau from HCT observations matches well with that obtained from the LAMOST spectra. We are unable to obtain a multi-epoch of observation for Anon 1 and HD 283782. In column 5, we have listed average values of EW(H) corresponding to each TTS, and errors associated with it indicate the range in EW(H). As the strength of H provides an indirect indication of accretion rate, a large range in EW(H) in CIDA 1, CW Tau, and J041411 suggests a large variation in accretion rate in these TTS.
@@
The H line strength has often been used to distinguish between CTTS and WTTS in the literature (White & Basri, 2003). WTTS show weak emission due to chromospheric activity, while comparatively strong emission from CTTS is produced due to accretion. However, there is no unique value of EW(H) that act as a dividing line for this classification (Barrado y Navascués & Martín, 2003; White & Basri, 2003). Barrado y Navascués & Martín (2003) provided EW(H) values as a function of spectral types, derived from the observed saturation limit for the chromospheric activity (blue line in Figure 5). Whereas, (White & Basri, 2003) provided maximum EW(H) values for WTTS for different ranges in spectral types (black dashed line in Figure 5). We have over-plotted average EW(H) values of our sample TTSs in Figure 5. The spectral types are estimated from photospheric temperature based on the correlation given in Pecaut & Mamajek (2013), listed in Table 3. Figure 5 shows that FM Tau, BS Tau, CW Tau, CIDA 1, FN Tau, and FO Tau have high values of EW(H) with respect to the expected values from WTTS with similar spectral types. These stars also show high UV excess, as discussed in the previous section, indicating that their classification as CTTS is robust. Except for BS Tau and FN Tau, all other sources discussed above show significant variation in the EW(H), suggesting a large variation in the accretion rate. In the next section, we estimate and discuss the accretion luminosity and mass accretion rate of these sources. We, however, need more followup observations for a detailed study of their variability.
@@
Anon 1, V773 Tau, and HD 283782 have equivalent widths less than that expected for their spectral types, which indicates that these TTS can be classified as WTTS. From the SEDs, we have also seen relatively low UV excess over photospheric emission. Both these scenario suggests that these are WTTSs. Figure 5 also shows that V836 Tau falls on the line separating CTTS from WTTS and the range in the EW(H) extends on both sides, making it difficult to classify. However, we notice a large amount of UV excess emission as shown in Figure 4, which is comparable to that found in other CTTS, suggesting that V836 Tau is a CTTS. We also notice that V836 Tau changes its EW from 8.88 to 17.6 in the two epochs of observation, which corresponds to 100% increase in the EW. These observations suggests that the star is going through a variable accretion rate. Therefore, based on both UV excess and EW(H), we suggest that V836 Tau is a slowly accreting CTTS with a variable accretion rate. Further monitoring observations of V836 Tau is required to understand its accretion properties. Our classification based on the spectroscopic observations is listed in the Table 2.
@@
Our classification of these sources matches well with the previous literature. From the literature, we find that FM Tau, CW Tau, FN Tau and FO Tau are classified as CTTS, while V773 Tau and HD283782 are classified as WTTS (Herbig & Bell, 1988; Kenyon et al., 1998; Valenti et al., 2003; Ingleby et al., 2013; McClure et al., 2013; Gómez de Castro et al., 2015), which agrees with our classification. BS Tau and V836 Tau are classified as WTTS based on the UV-NIR colour-colour relation (Gómez de Castro et al., 2015), contrary to our classification of CTTS for both sources. Based on HST FUV spectroscopic observations, V836 Tau is classified as slowly accreting CTTS (Ingleby et al., 2013). However, both UV excess and EW(H) values from our analysis indicate that BS Tau and V836 Tau are CTTS. We did not find any classification for CIDA 1, J0414, and Anon 1 in the literature. Thus from our analysis, we reconfirm that V836 Tau is CTTS and for the first time we report BS Tau, CIDA 1, and J0414 are CTTS and Anon 1 is WTTS. This study demonstrates the importance of multi-wavelength SED analysis and in particular, the simultaneous FUV and NUV observations to classify the TTS as classical or weak-line.

### v2

@@ 2.3 Spectral energy distribution
Out of 10 UVIT-detected TTSs candidates, only six are detected in both FUV and NUV bands, while the others are detected only in the NUV bands. Therefore, we performed two different approaches to fit the observed SEDs to estimate the stellar parameters. We made a two-component fit to the UV and optical part of the SEDs of the six TTSs detected in both FUV and NUV bands; we used a stellar photospheric model and a blackbody (which represents the excess over the photosphere emission due to accretion shocks) to fit the observed SEDs to determine the excess UV luminosity, its equivalent temperature, and stellar parameters of the stars. We did not include NIR regions for the SED fit as the excess emission in the NIR region comes from the disk. For the remaining four TTSs, we only fit the photospheric models of dwarf stars to the optical part of the observed SEDs to estimate stellar parameters.
To construct observed SEDs, we cross-matched our sample with Gaia-DR3 (Gaia Collaboration et al., 2021; Babusiaux et al., 2022), APASS DR9 (Henden et al., 2015), and PanSTARRS (Magnier et al., 2020) for the optical; and 2MASS (Skrutskie et al., 2006) for the NIR region. We also included GALEX GR6/7 data (Bianchi et al., 2017) for this analysis. We have used the virtual observatory (VO) tool, VOSA (VO SED Analyzer; Bayo et al., 2008) to generate SEDs by converting magnitudes into flux values for corresponding filters and then to fit the observed flux distribution with theoretical model spectra. We have obtained the distances from the Gaia-DR3 (Bailer-Jones et al., 2021). VOSA performs multiple iterations to get best-fitted spectra to observed flux distribution by varying Teff, log(g), metallicity, extinction, and scaling factor () values and gives the best-fitted parameters after performing a minimization. As these TTSs are in the solar neighbourhood, we consider that they all have similar metallicity as the Sun. We kept extinction as a free parameter to estimate its value from the best-fitted spectra. The scaling factor () is used to scale the model flux to match the observed flux and is defined as where is the radius of the star and D is the distance to the star. For the stellar photosphere, we have used the BT-Settl-CIFIST model (Allard et al., 2011). We have used log(g) values between 4 to 5 and the full range for Teff from 1200 to 7000 K. In the case of the black body, we used a range of temperatures from 5000 to 20000 K.
@@
where N is the number of photometric data points, n is the number of input-free parameters, is the observed flux, and is the model flux. Though is used to determine the quality of the fit, often values are found to be larger despite visual inspection suggesting them to be well-fitted SEDs. This large value arises when the photometric data points have very small observational flux errors (say 1% of observed flux). So, even if the model reproduces the observation apparently well, the deviation can be much higher than the reported observational error (increasing the value of ; Rebassa-Mansergas et al., 2021). To mitigate this, VOSA has introduced the parameter called the visual goodness of fit (Vgfb) by modifying the formula, where the error is considered to be at least 10% of its observed flux333http://svo2.cab.inta-csic.es/theory/vosa/helpw4.php?otype=star&action=help. The parameter Vgfb provided by VOSA with a value 15 is usually considered as a proxy for well-fitted SEDs (Rebassa-Mansergas et al., 2021). We considered the Vgfb parameter for the goodness of fit, however, we mentioned both and Vgfb parameters from the SEDs fitting process.
@@
In Figure 2, we have shown SEDs of the six TTSs (detected in both FUV and NUV) fitted with model spectra. The observed fluxes are shown as cyan and blue points, where cyan points fall in the UV and optical regions of the energy distribution which is included in the dual component fitting of model spectra. The blue points fall in the NIR region and are not included in the fit. The black and grey lines represent the best-fitted black-body and BT-Settl-CIFIST spectra, respectively, to the cyan points. The combined flux of these two model spectra is shown as the red dashed line. The red points indicate the combined flux in different bands. The overlap between red and cyan points indicates the goodness of the two-component fitting. We notice that in the case of HD283782, combined flux matches very well in the NIR region (not included in the fit) as well. Therefore, there is no NIR excess present in HD283782, which indicates that the TTS does not have hot/warm dust around it. The values of reduced , Vgfb, temperatures, logg, and extinction (AV) corresponding to the best-fitted spectra are noted in the legends. The names of the sources are mentioned at the top of each plot. V836 Tau and HD283782 were found to have high values (100) due to the above-mentioned reason of very small photometric error, however, smaller values of Vgfb (5) indicate that SEDs are fitted well by the model spectra.
@@
In Figure 3, we have shown SEDs of TTSs with no FUV detections. Due to the absence of FUV points, only the optical region of the observed SED is fitted with synthetic dwarf spectra (BT-Settl-CIFIST) to get the stellar parameters of these TTSs. Fitting two-component spectra to these sources with no FUV data points leads to the wrong estimation of parameters for the blackbody and stellar spectrum. We arrived at this conclusion after examining two-component fits to those six TTSs having both FUV and NUV, but FUV data points were not included in the fitting process and found that the best-fitted spectra were not able to produce the observed FUV flux. We have used Teff, , and AV as free parameters, and metallicity to be fixed at Solar value. Blue and cyan points in Figure 3 are observed SEDs, where only cyan points are included in the fitting process. The grey spectrum indicates the best-fitted model spectrum to the optical region of the observed SED. The red point indicated the expected flux due to the best-fitted model spectrum. The best-fitted parameters are mentioned in the legend. Except for the source J0414, all the other three sources have very high values, however, all four sources have Vgfb 15 suggesting the model spectra fitted well to the observed SEDs.
@@ 3 Results and Discussion
The SED analysis provides us with fundamental parameters of the TTSs, i.e., Teff, logg, radius, mass, and bolometric luminosity (Lbol) of the central stars. As mentioned on the VOSA website, the fitting process and the predicted flux are relatively less sensitive to . This poses a challenge to put direct constraints on the masses of the components since a slight change in can create a large difference in the estimated mass. Therefore, we do not include the estimations of mass and in our analysis. The values of other parameters are listed in Table 3. The luminosity of the black-body provides a direct measure of the accretion luminosity/chromospheric luminosity. In Table 3, we listed the value of bolometric luminosity enclosed within the black-body spectrum (Lacc) and black-body temperature (TBB). We have also estimated the spectral type of each TTS using Teff-spectral type relation (Pecaut & Mamajek, 2013) and listed in column 7 of Table 3. We find that TTSs have a Teff range from 3000 K to 4600 K and spectral types of K and M. The TBB ranges from 8500 K to 10700 K.
3.1 UV excess as an indicator of accretion
@@
We notice in Figure 2 that V773 Tau and HD283782 emit negligible amounts of excess emission in NUV bands above the photospheric emission compared to that found in the FUV bands. Whereas, in the other four sources we observe a large excess emission in both NUV and FUV bands. In Figure 3, we further notice that Anon1 has less excess NUV bands compared to other sources. The presence/absence of NUV excess appears to divide sources into two different categories.
@@
However, we need to quantitatively estimate the amount of excess emission in all the sources and compare them before giving any conclusion. To quantify the excess flux observed in the FUV and NUV regions, we define excess emission in a particular band as . To demonstrate UV excess, we used the N245M filter in the NUV and the F148W band in the FUV regions, as these are common bands for all the sources. We have plotted NUV and FUV excess as a function of effective temperature (Teff) in the upper panel of Figure 4 and as a function of blackbody temperature (TBB) in the lower panel. Teff and TBB are estimated from SED fitting. The sources detected in both FUV and NUV are marked as blue points, while the sources with only NUV detection are marked in orange. We can see that there are two groups of population, independent of their spectral types and blackbody temperature, separated by a black dashed line. The presence of high UV excess (103 in NUV and 107 in FUV) suggests that these seven sources (FM Tau, CW Tau, BS Tau, V836 Tau, CIDA1, FO Tau and J0414) could be still actively accreting, while V773 Tau, HD283782 and Anon1 are probable non-accreting WTTSs candidates with comparatively little excess in UV. Comparatively low UV excess with almost no NIR excess in HD283782 indicates that it is an example of disk-less WTTS. A more careful inspection suggests that stars with higher UV excess (or the probable CTTSs) also tend to have hotter TBB compared to the stars with lesser UV excess. However, we need more sources with simultaneous UV observations to confirm this aspect. We classify the stars emitting 103 (107) excess emission in NUV (FUV) as CTTSs, while the other group of stars are classified as WTTSs. Our classification based on the UV excess is listed in the Table 1. However, a strong UV flare might also cause excess UV emission and appear as CTTSs in Figure 4. Therefore, we require further evidence from spectroscopic observations before confirming their classification as CTTSs or WTTSs, as CTTSs are expected to show strong H emission compared to WTTSs.
| TTS | TBB | Teff | Radius | Lbol | Spectral Type | Lacc | ||
| Name | (K) | (K) | (R⊙) | (L⊙) | (L⊙) | (10-9 M⊙yr-1) | ||
@@ 3.2 H line emission as an accretion indicator
We analyzed spectra obtained from both the HFOSC/HCT and LAMOST for our sources. We estimated the equivalent width of H line emission (EW(H)) after subtracting continuum flux from the spectra. In Table 2, we have listed the EW(H) along with its measurement error for each source obtained from the multi-epoch observations with different telescopes (column 4). Names of the telescopes and the dates of observations are also mentioned in columns 2 and 3, respectively. We notice that EW(H) of FM Tau and V773 Tau from HCT observations matches well with that obtained from the LAMOST spectra. We are unable to obtain a multi-epoch of observation for Anon 1 and HD 283782. In column 5, we have listed average values of EW(H) corresponding to each TTS, and errors associated with it indicate the range in EW(H) from multi-epoch observations ( ). All the sources, except J0414, are found to have smaller measurement uncertainties in EW(H) compared to its range from multi-epoch observations. As the strength of H provides an indirect indication of accretion rate, a large range in EW(H) with relatively smaller observational errors in CIDA 1 and CW Tau suggests a large variation in accretion rate in these TTSs.
@@
The H line strength has often been used to distinguish between CTTSs and WTTSs in the literature (White & Basri, 2003). WTTSs show weak emission due to chromospheric activity, while comparatively strong emission from CTTSs is produced due to accretion. However, there is no unique value of EW(H) that acts as a dividing line for this classification (Barrado y Navascués & Martín, 2003; White & Basri, 2003). Barrado y Navascués & Martín (2003) provided EW(H) values as a function of spectral types, derived from the observed saturation limit for the chromospheric activity (blue line in Figure 5). Whereas, (White & Basri, 2003) provided maximum EW(H) values for WTTSs for different ranges in spectral types (black dashed line in Figure 5). We have over-plotted values of our sample TTSs in Figure 5. The spectral types are estimated from photospheric temperature based on the correlation given in Pecaut & Mamajek (2013), listed in Table 3. Figure 5 shows that FM Tau, BS Tau, CW Tau, CIDA 1, FO Tau and J0414 have high values of EW(H) with respect to the expected values from WTTSs with similar spectral types. These stars also show high UV excess, as discussed in the previous section, indicating that their classification as CTTSs is robust. Except for BS Tau, all other sources discussed above show significant variation in the EW(H), suggesting a large variation in the accretion rate. J0414 has a large uncertainty in the estimation of EW(H), so variation in J0414 is most likely not due to accretion variability. In the next section, we estimate and discuss the accretion luminosity and mass accretion rate of these sources. We, however, need more follow-up observations for a detailed study of their variability.
@@
Anon 1, V773 Tau, and HD 283782 have equivalent widths less than that expected for their spectral types, which indicates that these TTSs can be classified as WTTSs, marked as diamond points in Figure 5. From the SEDs, we have also seen relatively low UV excess over photospheric emission. Both these scenario suggests that these are WTTSs. Figure 5 also shows that V836 Tau falls on the line separating CTTSs from WTTSs and the range in the EW(H) extends on both sides, making it difficult to classify. However, we notice a large amount of UV excess emission as shown in Figure 4, which is comparable to that found in other CTTSs, suggesting that V836 Tau is a CTTS. We also notice that V836 Tau changes its EW from 8.88 to 17.6 in the two epochs of observation, which corresponds to 100% increase in the EW. These observations suggest that the star is going through a variable accretion rate. Therefore, based on both UV excess and EW(H), we suggest that V836 Tau is a slowly accreting CTTS with a variable accretion rate. Further monitoring observations of V836 Tau are required to understand its accretion properties. Our classification based on the spectroscopic observations is listed in the Table 2.
@@
Our classification of these sources matches well with the previous literature. From the literature, we find that FM Tau, CW Tau and FO Tau are classified as CTTSs, while V773 Tau and HD283782 are classified as WTTSs (Herbig & Bell, 1988; Kenyon et al., 1998; Valenti et al., 2003; Ingleby et al., 2013; McClure et al., 2013; Gómez de Castro et al., 2015), which agrees with our classification. BS Tau and V836 Tau are classified as WTTSs based on the UV-NIR color-color relation (Gómez de Castro et al., 2015), contrary to our classification of CTTSs for both sources. Based on HST FUV spectroscopic observations, V836 Tau is classified as a slowly accreting CTTS (Ingleby et al., 2013). However, both UV excess and EW(H) values from our analysis indicate that BS Tau and V836 Tau are CTTSs. We did not find any classification for CIDA 1, J0414, and Anon 1 in the literature. Thus from our analysis, we reconfirm that V836 Tau is a CTTS and for the first time we report BS Tau, CIDA 1, and J0414 are CTTSs and Anon 1 is a WTTS. This study demonstrates the importance of multi-wavelength SED analysis and in particular, the simultaneous FUV and NUV observations to classify the TTSs as classical or weak-line.